A child is prescribed amoxicillin at a dose of 45 mg/kg daily to be given in three divided doses at 8 hour intervals. If the child weighs 12 kg, what volume of amoxicillin suspension containing 100 mg/5ml should be given every 8 hours? BC
Answer #1

ANSWER

Amount of amoxicillin given every 8 hours.

         Formula

              Amount want to give (x) = Dose ordered(D) *Quantity(Q) / Have (H).

         Dose ordered = 45 mg/kg/day.

                Weight of the child = 12 kg

   so

Dose ordered (D) = 45 * 12

                                                     = 540 mg / day.

                                                          = 540 / 3

                                                          = 180 mg / 8hr

                          Quantity(Q) = 5 ml.

                                        Have (H) = 100 mg.

               Amount should give 8 hrly

                            As per formula

                                          X = D * Q / H

                                             = 180 * 5 / 100

                                             = 900 / 100

                                             = 9 ml (q8h)